Top Physics at Hadron

Colliders

Pekka K. Sinervo University of Toronto / IPP

Lecture # 1

1. Introduction to Top Quark Physics 2. Production in Hadron Collisions

3. Top Decays and Detector Requirements 4. D0 / and CDF Experiments

5. Multijet Decays

6. Dilepton Searches

Lecture # 2

1. Lepton + Jet Searches

2. B Quark Tagging

3. Results from Run 1A/1B

4. Cross section and mass estimates

5. Outlook & Conclusions

Lake Louise Winter Institute February 1995

Introduction: Why a Top Quark?

The Standard Model of Elementary Particle Physics

• Three “generations” of quarks and leptons

( \begin{matrix}u \\ d \end{matrix} ) ( \begin{matrix}c \\ s \end{matrix} ) ( \begin{matrix}t \\ b \end{matrix} )
( \begin{matrix}e \\ ν_{e} \end{matrix} ) ( \begin{matrix}μ \\ ν_{μ} \end{matrix} ) ( \begin{matrix}τ \\ ν_{τ} \end{matrix} )

“Isospin” doublets that dictate particle interactions

• Forces carried by bosons

Electromagnetic force  γ
Strong force  eight gluons
Weak force  W ±, Z∘ bosons
Gravity  gravitons

SM Predictions for Top Quarks

The asymmetry in e^{+} e^{−} → bb̄ scattering implies that b is part of SU(2) doublet

• top quark has standard SU(2) couplings

• CKM matrix predicts dominate decay t → W + b

Corrections to W^{±} and Z^{∘} masses depend on heaviests fermions

• LEP, SLD measurements on the Z resonance require

Mtop = 178 ± 10{+ 18 \atop −19} GeV/c2

D. Schaile, CERN-PPE/94-162

Top Quark Phenomenology

Top Quarks decay via the weak force

• Predict 100% t → W + b

• Lifetime depends on mass

• For heavy top (Mtop > ∼ 100 − 120 GeV/c2), decays before it hadronises

Sensitive to some classes of “new physics”

• Higgs couples strongly to top

• Could decay into a SUSY H+

t → H+ b

|_{⟶}τ+ ντ

• Top quark can be sensitive to SM details – Vts CKM matrix elements

– Constraints on other SM parameters (eg. ALR)

What Do We Already Know About Top?

Searches at e^{+} e^{−} Colliders have ruled out “light” top

• data from LEP requires Mtop > 46 GeV/c2 at 95% CL • Independent of how top decays

Width of W ± boson rules out W + → tb̄ • Mtop > 62 GeV/c2 at 95% CL

Searches at pp̄ colliders at CERN and Fermilab

• searched for a “Standard Model” top

• Mtop > 131 GeV/c2 at 95% CL (D0 Collab).

Production of Top in Hadron Collisions

Top quark production is mediated by qq̄, gg → tt̄

qq̄ → tb̄X
qq̄ → W^{+} → tb̄

Because top is heavy, need large x for initial partons

• Implies that cross section falls with mass

• t t system approximately at rest – < PT >∼ Mtop/2

Higher order corrections to σ_{t t} have been calculated

Top Decays and Detection Requirements

SM top quark decays via the weak force

t → W + b

|_{⟶}ud̄ or cs̄ (2/3)

|_{⟶}e+ νe or μ+ νμ or τ+ ντ (1/3)

In proton-antiproton collisions,

This results in several distinct decay signatures

• Six jet final states

• Lepton + Jet events with at least one b quark

• Dilepton events with two, energetic e± and μ∓

• Decay modes with τ’s hard!

Detector Requirements

Detection of these decay modes requires

• Good jet reconstruction (energy and direction)

• Good lepton ID capability (acceptance and efficiency)

• Missing energy detection (for neutrinos)

Production/decay properties imply that

• Production is “central” (especially for high mass) – Most likely initial state is where x1 ∼ x2 – Top quarks are produced with modest PT

⇒ daughters are isotropic in lab frame

Detector requirements boil down to

1. Coverage over |η| < ∼ 2 − 2.5 desirable

2. Good efficiency for partons with PT ≥ 15 − 20 GeV/c 3. Good lepton ID with high efficiency

Can see this from t t MC with M_{top} = 160 in dilepton channel

Background Considerations

The backgrounds are dominated by QCD processes

• Multi-jet final states direct background to 6-jet channel • Jets faking leptons potential background to lepton ID • Many SM EWK physics processes

– Drell-Yan production in e+ e− and μ+ μ− modes

– Z∘ production

– semi-leptonic bottom and charm quark decays

– QCD W + jet processes

– Di-vector boson final states: W + W −, W + Z∘

Any of these could compromise one of the searches

• Lepton ID must have extremely good jet-rejection

• Kinematics of background events must be used

– For example, can use PT and lepton isolation requirements to reduce and characterise b, c quark production

• Bottom line:

Background Studies Extremely Important

The Fermilab Collider Experiments

CDF and D0 / designed to study p p interactions

• Collisions with √s = 1.8 TeV created by Tevatron Collider

• Sophisticated general-purpose detectors operated by big collaborations

Main detector features:

• Hermetic Calorimetry – D0 / has state-of-the-art Ur/Li Ar calorimeter

– CDF has more traditional Pb and Fe calorimeter ∗ Scintillator readout for central region |η| < 1.1

∗ Gas proportional chamber readout for plug/forward – Coverage over |η| < ∼ 4

– Good hermiticity for E̸T measurement

– Provides electron ID and jet detection

Detectors (continued)

• Muon chambers outside calorimeter – identify charge tracks that pass through calorimeter

– D0 / has toroidal magnets to measure muon PT

• Central Tracking Systems

– CDF has three-tiered system in 1.4 T field ∗ Central Track Chamber (CTC) for 28 < R < 130 cm ∗ Vertex Chambers (VTX) in 10 < R < 23 cm

∗ Silicon Vertex Detector (SVX) in 3 < R < 12 cm

– D0 / has nonmagnetic central tracker

∗ Also incorporates a Transition Radiation Detector (TRD) for electron ID

By design, detectors are complementary

• D0 / emphasises calorimetery and muon ID

• CDF emphasises tracking and electron ID

• Reality is that both have strengths and weaknesses

Detector Limitations

Both detectors have some limitations. CDF’s dirty laundry”

• Muon system – not enough absorber

– coverage somewhat fragmented

• Calorimetry – Not totally hermetic

– Region |η| > 3.6 not well understood

D0 / ’s laundry list:

• Trigger system still being shaken down

• Lack of magnetic field reduces redundancy – Calibration more difficult

– Not able to detect secondary vertices

Triggering Issues

Detectors use 3-level trigger system to identify events in real-time

• Interaction rate of (50 × 10−27 cm 2) ×(1.5 × 1031 cm −2s−1) = 7.5 × 105 Hz has to be reduced to of order 1-10 Hz

• Most challenging aspect of experiment – Depends critically on all subsystems

– Need efficiency > 75% to be competitive

Triggering for “heavy” top is relatively easy

• Tag high PT lepton(s)

• Missing energy requirements

• Multijet topologies

• Trigger efficiencies ≥ 90 − 95% in most cases

1992-95 Collider Run

Tevatron Collider has been running since 1992

• Run 1A (1992-93) provided ∼ 20 pb−1

– CDF scanned ∼ 1 × 1012 collisions in 19.3 pb−1 ∗ Collected 16 × 106 events on tape

– D0 / collected 13.5 pb−1

• Run 1B (1994-95) has produced another ∼ 55 pb−1 – D0 / has updated results on top search (+22 pb−1)

– CDF is currently preparing preliminary results – Run expected to continue till end of 1995

• Goal is to have in excess of 100 pb−1

By far the largest and most sensitive dataset to top

• For a top quark with a mass of 160 GeV/c2

– σt t = 8.2 pb ⇒ 160 produced t t events

– expect to observe 1.3 dilepton events

– expect to observe 3.5 lepton + jet events

Dilepton Top Quark Search

t t decaying into 2 leptons create unique signature

• Two charged leptons from W ± decay

• Imbalance in total energy due to undetected neutrinos • Two jets from bottom quarks

D0 / and CDF performed a search requiring

• Two clean, high PT lepton candidates

• Missing transverse energy E̸T > 25 GeV

• Require two jets with ET > 15 GeV and |η| < 2.0 • Reject specific backgrounds (e.g. Z∘)

Significant dilepton background is Drell-Yan production

• E̸T signature created by mismeasured jet

• Harden E̸T cut if correlated with jet

CDF Results for Dilepton Channel

Have measured the rate of events from non-top sources

• Have used data to measure these rates

• Cross-checked using Z∘ → e+ e− and Z∘ → μ+ μ−

• Consistent with theoretical calculations

Background  eμ  ee, μμ
Z → τ+ τ−  0.07 ± 0.02  0.06 ± 0.02
One or both leptons fake  0.03 ± 0.03  0.04 ± 0.03
bb̄  0.04 ± 0.03  0.05 ± 0.03
W W, W Z and ZZ  0.10 ± 0.04  0.06 ± 0.02
Drell-Yan  0.10{+0.23 \atop −0.08}
Total  0.24 ± 0.06  0.31{+0.24 \atop −0.10}

CDF observes 2 eμ events

• Expect 0.56{+0.25 \atop −0.13} background events

D0 / Results for Dilepton Channel

D0 / applies slightly different P_{T} cuts on leptons

• Goal is to optimise signal to noise

• Lower lepton thresholds for eμ channel

• Jet cuts reject larger background fraction

They observe two events in their 1A and 1B data

eμ  e+ e−  μ+ μ−
Data (1A)  1  0  0
Data (1B)  1  0  0
Backgrounds  0.7 ± 0.1 0.5 ± 0.1 0.8 ± 0.1
Mtop = 140 GeV/c2  2.3  1.2  0.6
Mtop = 160 GeV/c2  1.2  0.7  0.3
Mtop = 180 GeV/c2  0.7  0.4  0.1

Observe 2 events and expect 2.0 ± 0.2 background

• Not very sensitive to top

Summary of Dilepton Results

CDF/D0 / observe four eμ events

• For CDF, result appears more significant – Both events have electron and muon candidates – One has signature of b quark jet

• For D0 / , observation consistent with background

Expected rate from top is small and depends on mass

Mtop (GeV/c2) CDF events  D0 / events
140  2.2  3.5
160  1.3  1.9
180  0.7  1.1

Overall significance is low

• Probability 0.56 events fluctuate to ≥ 2 events is 12%

• D0 / result totally consistent with background ⇒ Not statistically significant

Mass Limits from Dilepton Data

CDF uses dilepton data to exclude “low mass” top

• In this case, CDF removes jet requirement – Jet efficiency less certain for Mtop ∼ 100 − 120 GeV/c2 – No additional events are added to sample

• Add 3.7 pb−1 from 1988-89 data run (no events)

• Backgrounds increase to 3.0 ± 0.6 events

• Assume 2 events are top and set limit on σtop

D0 / has chosen not to update its lower mass limit Dilepton acceptance as a function of top mass

22

So What About Multijets?

t t → 6j faces huge QCD multijet background

• QCD cross section is over 4 orders of magnitude larger • Have been some studies

– Can use kinematic cuts to get of order 103 rejection – Use b quark tagging techniques to get ×10

• Still left with signal-to-background of 1-10 to 1-50! ⇒ most people have given up!